# Are Biological Systems More Intelligent Than Artificial Intelligence?

### v3

@@ Abstract
Is a biological self-organising system more ‘intelligent’ than an artificial intelligence? If so, why? We frame intelligence as adaptability, and explore this question using a mathematical formalism of enactive causal learning. We extend it to formalise the multilayer, multiscale, bottom-up distributed computational architecture of biological self-organisation. We then show that this architecture allows for more efficient adaptation than the static top-down interpreters typically used in computers. To put it provocatively, biology is more intelligent because cells adapt to provide a helpful inductive bias, and static interpreters do not. We call this multilayer-causal-learning. However it inherits a flaw of biological self-organisation. Cells become cancerous when isolated from the collective informational structure, reverting to primitive transcriptional behaviour. We show that, in the context of our formalism, failure states like cancer occur when systems are too tightly constrained by the abstraction layer in which they exist. This suggests control should be distributed (bottom-up rather than top-down) to ensure graceful degradation. We speculate about what this implies for systems in general, from machine learning hardware to human organisational and economic systems. Our result shows how we can design more robust systems and, though theoretical in nature, it lays a foundation for future empirical research.
@@ 1 Introduction
Biological systems self-organise, meaning their parts and the environment interact to produce spatiotemporal order [1, 2, 3, 4, 5, 6, 7]. They interact with their environment to achieve homeostatic and reproductive goals. Some proactively learn [8], storing information about the past in order to anticipate future needs. To make accurate predictions in environments of sufficient complexity, a system must be able to learn cause and effect [9, 10, 11]. For our purposes we consider the ‘intelligence’ of a system to be a measure of the resources it needs to correctly identify cause and effect [12, 13, 14, 15, 16] [^1] There are many definitions of intelligence, the merits of which are beyond the scope of this paper. Ours is primarily informed by [12], which is adapted from Legg and Hutter’s earlier definition of intelligence as ‘the ability to satisfy goals in a wide range of environments’. The more efficient the learning system, the wider the range of environments in which it can satisfy goals. It is not possession of skills, but an ability to acquire skills with limited resources. Being an embodied formalism, this also accounts of energy and other resources consumed.. It is a measure of how efficiently one adapts [^2] Primarily sample and energy efficiency. A ’sample’ is information used to learn. When a teacher provides an example of a math problem being solved, they are providing a sample. If student A requires 4 samples to understand a concept, and student B requires 5 samples, then A is more sample efficient than B..
Artificial intelligence (AI) is the study of adaptive machines, and machine learning (ML) of those that learn [17, 18, 19, 20]. Contemporary approaches to machine learning are notoriously inefficient [21]. They need vast quantities of data and energy learn tasks most humans can pick up after one or two examples. Even organoids, small collectives of cells grown in a lab, appear to be more ’intelligent’ in this sense than contemporary machine learning [22]. Why is this the case? Are biological systems more intelligent, or is something else going on?
@@
Some suggest that the relative inefficiency of ML may be down to the fact that biological systems are active inference systems, and machine learning systems are typically not [23, 6, 24]. Active inference is much more efficient [25, 26, 27], but it may only be part of the picture. As of 2024 active inference systems [^3] In particular the models submitted by Jack Cole and his employer, Minds AI. applied in the context of the ARC benchmark [14] [^4] For additional information, see https://arcprize.org for sample efficient adaptation still under-perform humans by a substantial margin [^5] Though it should also be noted that the Minds AI active inference system is one of the leading ARC submission for 2024, excluding OpenAI’s o3 which may or may not be an active inference system (the details have not been released at this time), and which is disqualified from ARC because it does not meet the computational efficiency requirements.. Others hold that the reason biological systems appear to be more intelligent is because of an ‘inductive bias’ [14]. This means the system embodies hard-coded foreknowledge of the environment, which predisposes it to efficient adaptation. If an inductive bias were the only reason for the discrepancy between biological self-organisation and machine learning, that would suggest biological systems are not more intelligent than artificial intelligence, but more knowledgeable.
To reiterate, we seek to explain why biological self organisation seems to be more sample (and energy) efficient than contemporary machine learning. Is there something biological systems do that allows them to be more ’intelligent’ than artificial intelligence in its contemporary form? We suggest the answer lies in how biological self organisation distributes control at a different scales, and levels of abstraction. To do so we must first give some background on causal learning which informs our hypothesis.
1.1 Causality
Cause and effect are typically represented as an acyclic graph.
@@ 1.1 Causality
Complex environments often complicate the ability to identify causes. For example, confounded events may seem to cause one another when in fact they are merely correlated. This can lead one to draw erroneous conclusions about cause and effect.
not the causeeffectconfounder
@@
By actively intervening in the environment a system can experiment to isolate causes. To do so, a system must be able to differentiate between events it passively observes and events it has caused [10, 9, 28]. Then, it can construct a causal graph. There are two ways to do that. One option would be to assume the nodes (variables, actions and other abstract objects and properties), and then learn the causal relation between them. This works well in theory for artificial intelligence [29], but requires foreknowledge of objects and properties between which causal relations can be established. The alternative is to assume the causal relations (the edges between nodes in the graph), and then learn the variables (the abstract concepts). This needs only attraction and repulsion from physical states [30], otherwise known as valence. We take this latter approach.
@@
We can frame causal learning in computational terms. Learning nodes in a causal graph is like figuring out which computer program was used to generate a sequence. Assume we have fed a set of inputs into a computer, and for each input the computer responded with a particular output. The computer is running a ’policy’ program which maps inputs to outputs [^6] We’ll call all such programs policies for the sake of this paper. This should not be confused with the more specific notion of a policy from reinforcement learning [19].. That policy is the cause of the input-output history we’ve observed so far. We want to know the policy that caused historical outputs so that we can use it to predict future outputs. The problem is that more than one policy exists which could have produced the history we’ve observed, and each of those policies produce different outputs when we give the computer inputs it hasn’t encountered before. From a naive Bayesian perspective, all of them are equally likely to have been the true underlying cause. If we wish to know which policy actually caused the outputs, we need way to narrow down the possibilities even further.
@@
There are two ways to compare policies that look the same with respect to historical data, but differ with respect to possible future data [31]. One way is to can compare their form, and choose programs which are simpler. However, in an interactive setting simplicity is an unreliable indicator of plausibility [32, 31]. The alternative is to compare policies by their function, in terms of how they behave given inputs for which we do not have historical data. Adaptive systems frequently satisfy the same functional constraints by different means [33], so we can compare different sorts by the constraints they satisfy. Each policy implies a constraint on possible outputs given a set of possible inputs. The weaker that constraint is, the greater the variety of future behaviours it permits. Previous research showed that if two policies are identical with respect to historical data, then the policy most likely to have generated historical data is the weaker of the two [16]. In theoretical terms the policy most likely to have caused behaviour is the weakest, and in experiments a system which preferred weaker policies outperformed one that preferred the simpler policies by $110-500\%$ [34]. This has been used to establish local upper bounds on embodied ‘intelligence’ [12].
@@
ML systems do not typically optimise for either weaker constraints on function, or simplicity of form. This at least partially explains their relative sample inefficiency [35, 34]. However, even when ML systems are modified to optimise for weakness, their capacity to do so limited by the “abstraction layer”. An abstraction layer is like a causal isolator [36] between a software “mind” and its environment. For example, the Python programming language is interpreted by the Python interpreter. In the end, Python gets translated into machine code, which hardware like an Apple M2 system on a chip then interprets to physical effect based on the ARM instruction set architecture. An abstraction layer is just physical embodiment by another name. Different abstraction layers permit different upper bounds on policy weakness, and thus different upper bounds on embodied ’intelligence’ [12]. In other words, physical embodiment imposes severe limitations, and the behaviour of software is determined entirely by the hardware and software abstraction layers in which it runs. Intelligence is primarily a matter hardware, not software. In enactivist terms, an intelligent choice of abstraction layer is one that reduces the “big world” of all information to a “small world” of relevant information [37, 38, 39].
@@
Put another way an embodied policy is a physical structure, like a tool. A tool can serve more than one purpose. For example, a hammer is both a weapon and a paperweight. A more specific tool completes fewer possible tasks, while a tool that implies a weaker constraint completes more possible tasks [21]. By preferring weaker policies, one avoids unnecessary assumptions and so adapts as efficiently as possible, meaning the weakest policies complete a ‘wider variety of tasks’ than all others [^7] As an informal epistemological razor, this means “explanations should be no more specific than necessary”. Intuitively, this is a similar idea to the principle of maximum entropy [40, 41], but it accounts for the subjectivity of complexity and the teleological nature of language [31]. For further intuition, a corollary to the idea that a weaker policy is more likely to be the cause of data is the inventors paradox, or equivalently the contravariance principle proposed by Cao and Yamins [42]. They argue tasks that more tightly constrain the space of correct policies make better training sets. A larger, more varied set of inputs and outputs implies a smaller set of correct policies. That makes it easier to identify the policy that generated the inputs and outputs. This supports the claim that weaker policies are more likely to be the cause of behaviour, because a weaker policy is more likely to be a correct policy for tasks that more tightly constrain the set of correct policies. See proofs 1 and 2 for more details.. To complete a ‘wider variety of tasks’ in the embodied sense, a tool must be both more sample and more energy efficient, and so the weakest policy is the most efficient in both senses.
That biological self-organising systems adapt efficiently suggests they are able to embody weaker policies than today’s ML systems. They attain higher local upper bounds on embodied intelligence [12]. This suggests they have a means of optimising not just the choice of policy, but the embodied abstraction layer in which that policy exists.

### v4

@@ Abstract
Are biological self-organising systems more ‘intelligent’ than artificial intelligence? If so, why? We frame intelligence as adaptability, and explore this question using a mathematical formalism of causal learning. We compare systems by how they delegate control, illustrating how this applies with examples of computational, biological, human organisational and economic systems. We formally show the scale-free, dynamic, bottom-up architecture of biological self-organisation allows for more efficient adaptation than the static top-down architecture typical of computers, because adaptation can take place at lower levels of abstraction. Artificial intelligence rests on a static, human-engineered ‘stack’. It only adapts at high levels of abstraction. To put it provocatively, a static computational stack is like an inflexible bureaucracy. Biology is more ‘intelligent’ because it delegates adaptation down the stack. We call this multilayer-causal-learning. It inherits a flaw of biological systems. Cells become cancerous when isolated from the collective informational structure, reverting to primitive transcriptional behaviour. We show states analogous to cancer occur when collectives are too tightly constrained. To adapt to adverse conditions control should be delegated to the greatest extent, like the doctrine of mission-command. Our result shows how to design more robust systems and lays a mathematical foundation for future empirical research.
@@ 1 Introduction
Biological systems self-organise, meaning their parts and the environment interact to produce spatiotemporal order [1, 2, 3, 4, 5, 6, 7]. They pursue homeostatic and reproductive goals. Some proactively learn [8], storing information about the past in order to anticipate future needs. To make accurate predictions in complex environments, a system must learn cause and effect [9, 10, 11]. We consider a system’s ‘intelligence’ to be a measure of how efficiently it does this [12, 13, 14, 15, 16] [^1] There are many definitions of intelligence, the merits of which are beyond the scope of this paper. Ours is primarily informed by [12], which frames intelligence as the ability to complete a wide range of tasks. This is an adapted from Legg and Hutter’s earlier definition of intelligence as ‘the ability to satisfy goals in a wide range of environments’. Considering a ‘wide range of tasks’ accounts for embodiment, as well as environment and, being an embodied formalism, this also accounts of energy and other resources consumed. Intelligence by this definition is not possession of skills, but an ability to acquire skills with limited resources.. Put another way, ‘intelligence’ is a measure of adaptability [^2] Primarily sample and energy efficiency. A ‘sample’ is information used to learn. When a teacher provides an example of a math problem being solved, they are providing a sample. If student A requires 4 samples to understand a concept, and student B requires 5 samples, then A is more sample efficient than B. Increased sample and energy efficiency means an increase in the total number of tasks one can complete, and thus adaptability..
Artificial intelligence (AI) is the study of adaptive machines, and machine learning (ML) is the study of machines that learn [17, 18, 19, 20]. Contemporary approaches to machine learning are notoriously inefficient [21]. They need vast quantities of data and energy to learn tasks a human can pick up after one or two examples. Even organoids [^3] By organoid we mean a collective of cells grown in lab, placed in a computational interface and life support system, and then ‘trained’ to communicate with conventional computing technology. appear more ‘intelligent’ in this sense [22]. For example, an organoid outperformed conventional reinforcement learning systems at the game Pong, adapting with greater sample efficiency. That organoid was a highly tuned hybrid system whose behaviour was shaped by human-imposed boundary conditions. This is similar to how an acellular slime mold can be made to ‘solve’ shortest-path problems [23]. These sorts of results are highly dependent upon the human imposed boundary conditions. Nevertheless, they raise an interesting question. Are biological self-organising systems more intelligent than artificial intelligence? If so, why?
@@
Some suggest that a key difference between biological systems and contemporary artificial intelligence is that the former are active inference systems, while the latter are typically not [24, 6, 25]. It is true that active inference can improve efficiency [26, 27, 28]. Active inference has been applied [^4] In particular the models submitted by Jack Cole and his employer, Minds AI. in the context of the ARC-AGI benchmark [14] [^5] For additional information, see https://arcprize.org. for sample efficient adaptation. Active inference systems perform comparatively well, but still under-perform humans [^6] Though it should also be noted that the Minds AI active inference system is one of the leading ARC submission for 2024, excluding OpenAI’s o3 which may or may not be an active inference system (the details have not been released at this time), and in any case o3 is extremely inefficient.. Hence there may be more to the picture.
Active inference aside, some hold biological systems are not more intelligent. They only seem to be because of a ‘strong inductive bias’ [14]. This means the system embodies hard-coded foreknowledge of the environment, which predisposes it to adapt. To whatever extent this is true, it merely substitutes the question ‘what is intelligence?’ for ‘what is the inductive bias?’.
To reiterate, we want to know why biological self-organisation seem more adaptive [^7] Sample and energy efficient. than contemporary AI. Is there something biological systems do that allows them to be more ‘intelligent’? Could there be more to the picture than active inference and an inductive bias?
We suggest there is. A notable feature of biological self-organisation is the multiscale competency architecture (MCA). Adaptation takes place at every scale, from cells to organs to organisms to entire ecosystems [29]. It is ‘scale-free’. Similar dynamics tend to repeat at different scales. Control is delegated. Cells are a “soft” material that has agency [30, 31]. A cell has a perceptual field. A cell can co-operate with other cells in this perceptual field. This allows them to form multicellular networks that have larger a perceptual field than any individual cell, to maintain coherent higher level “identity” like that of an organism [32]. Each cell has its own goals and is intelligent in the sense of being able to reach the same goal by different means [33, 34]. However each cell can also be part a collective of cells that in turn pursues its own collective goals. Smaller collectives can form larger collectives, and the pattern repeats at ever larger scales. Collectives are nested [35, 36]. Each “thinks” in the enactive sense, interacting with its environment [37, 38, 39]. Biological collectives can exhibit high level goal directed behaviour far beyond the capability of the parts of which they’re formed [40].
Figure 1: The Barnsley Fern is an iterated function system fractal that exhibits self similarity [41]. Notice how the same pattern repeats at different scales. This gives some visual intuition as to what ‘scale-free’ suggests, as we can see the same dynamics play out again and again at different scales.
Our somewhat provocative claim is that biological self-organising systems are inherently more intelligent than contemporary implementations of artificial intelligence. We will argue this is because biological self-organisation delegates control to the lowest possible level of abstraction. We also present three proofs. The first two show that adaptability depends on the ability to embody “weak” constraints. The third proves the ability to embody weak constraints at higher levels of abstraction depends on the ability to embody weak constraints at lower levels. Our results, though purely theoretical, lay a foundation for future empirical research.
2 Abstraction and Delegation of Systems
Here we will briefly explain what we mean by abstraction and delegation. These are dimensions along which we may plot any system. For intuition, we will give examples of limit cases for computational, biological, human organisational and economic systems.
2.1 Abstraction
Abstraction is about how a system produces emergent, increasingly abstract behaviours. Highly abstract behaviours are a consequence of simpler goal directed behaviour at lower levels of abstraction. For example a collective of cells can enact an organ, and a system of organs can enact an organism [32] which enacts even more abstract behaviours like language or tools. Each of these is an ‘abstraction layer’ [12]. Abstraction layers exist in a stack (which we represent as graph). Intuitively, this is a ‘chain’ of cause and effect where each link in the chain (an abstraction layer) is the goal directed behaviour of the link (layer) that preceded it.
$cells$$organs$$organism$
When we say we examine a system at a lower level of abstraction, we mean we look at a node lower in the stack (or earlier in the chain). In the case of biological systems lower levels of abstraction are formed of smaller or simpler parts [42], so looking at a lower level of abstraction is like ‘zooming in’ to look at a smaller scale. Conversely, rather than looking at low levels of abstraction like cells we can shift the level of abstraction to higher in the stack, and look at individual humans and how they form collectives.
$soldiers$$squads$$platoon$
A conventional computing system is a stack of abstraction layers. For example a predictive model typically relies on a machine learning library, which runs on CUDA, which runs on an Nvidia GPU.
$GPU$$CUDA$$library$$model$
2.2 Delegated Control
Unlike biological systems a lower level of abstraction in a computer is not necessarily made up of smaller parts that form a collective. Looking at a lower level of abstraction is not like ‘zooming in’. For example, Python code is run by an interpreter, which can run on a computer. That interpreter is not a collective. It is not made up of parts capable of functioning independently.
Delegating work is not the same thing as delegating control. For example, work can be delegated to GPUs, but if we remove a GPU from the computer it will not then pursue goals independently. It will sit, inert until once again integrated into a larger system. It requires no incentives. Control is imposed top-down, from higher to lower levels of abstraction.
Conversely in a biological system each cell is independently capable of goal directed behaviour [31], assuming it remains alive. Control in a complex biological system begins bottom-up, before top-down. Control then propagates in both directions. Each and every part of a collective acts on its own, and those actions imply a ‘policy’. Each individual policy determines how that part interacts with the other parts, or more to the point their policies. These interactions cause the system as a whole to act, which implies a collective policy. If collective and individual policies align in service of homeostasis, then the system can persist. It has a collective identity or goal. That identity can constrain parts of the collective. For example, we might say the level of abstraction is cells, and then describe an organ as a collective policy of those cells. Likewise, we might look at the level of organs, and describe a human as a collective policy of those organs. Or we could describe the human as the collective policy of cells, and see the organs as a sort of middle management layer.
Figure 2: Illustration of a decentralised system (left) vs a centralised system (right). Notice that if a central node from the system on the right is deleted, then large portions of the graph become disconnected. If a node is deleted from the left graph, then there would still be a path between all remaining nodes. The delegation of control necessitates decentralisation.
A good example to understand the limit cases for delegation is an economy. In an ideal centrally planned economy control is concentrated at the highest level of abstraction. We can think of the people in the economy as a lower level of abstraction. Those people can adapt, but are not really allowed to because all decisions are imposed top-down. On the other extreme control is delegated to parts at the lowest level of abstraction, and conditions are such that control remains decentralised (i.e. there is some method of redistribution such as war or technological innovation).
When we choose a level of abstraction, we choose a smallest unit of delegation. For example if the level of abstraction is humans, then we don’t worry about how control is delegated within a human. We just concern ourselves with what systems can be formed of humans. Like an economy. Of course, there are limits on what sort of physical forms can store or process information in service of a goal [43]. This means there are limits to how much control can be delegated.
Figure 3: Examples of limit cases for delegation at different levels of abstraction.
2.3 Reasons to Suspect Delegation Facilitates Adaptation
We will now give a brief explanation of why we believe delegation of control is relevant to adaptation. All else being equal the lower the level of abstraction at which the system can be adapted, the more it can be specialised. For example a program written in C can be far more efficient than a program written in Python [44]. This is because the Python program has to be interpreted by the Python interpreter, which is a program written in C. By programming in C we simply cut out an abstraction layer. This gives the programmer more leeway to optimise the system, or make mistakes. Likewise another way to make computers more efficient is to use a field programmable gate array (FPGA) to write a program at the hardware level, instead of compiling that program into machine code and running it on a one-size-fits-all instruction set architecture like x86 [45].
With each successive abstraction layer introduced, one is further removed from the ground truth to which one must adapt. For example software is nothing more than a state of hardware [12]. It is a label we ascribe to the behaviour of hardware. It is a useful simplification that lets us consider what we want hardware to do in the abstract, but working at such a high level of abstraction can cause problems. Proposals for software ‘intelligence’ have been undermined by the fact that any claims regarding the behaviour of software are contingent upon hardware. The software doesn’t interact with the environment, the hardware does. This is because the ability to adapt is related to perceived complexity, which is entirely determined by the abstraction layer through which an agent interacts with its environment [46]. Consequently an ‘objective’ measure of artificial intelligence [47, 12] must be ‘enactive’ [37] in that it hinges not just software but hardware and the local environment as well. To understand the difference between biological systems and artificial intelligence, we must account for the level of abstraction at which an adaptive process takes place.
In distributed computational systems, like a drone swarm or network, delegating control can facilitate what is called ‘graceful degradation’ [48]. For example, to control a drone swarm we could single out one drone to be the ‘leader’, and have it run a program that co-ordinates the other drones. However that would make the swarm vulnerable. If the leader drone were damaged, the swarm would de-cohere. It is a single point of failure. In contrast if we decentralise control, delegating to the level of individual drones, the swarm will remain coherent regardless of how many drones are disabled. In other words, the swarm ‘adapts’ to adverse conditions.
The advantages of delegating control are found in human organisations as well. For example, in a military context there is a concept called ‘mission command’ [49]. A culture of mission command empowers units to adapt the situation at hand, by granting them leeway to bend rules and orders in service of the overall mission. It loosens the constraints under which units must operate, delegating control. Officers at every level are trained to take the initiative and act independently. A well documented example of the effectiveness of mission command in facilitating adaptation is that of Nordbat 2, a joint peacekeeping force deployed to Bosnia in the early 90s. Mission command allowed individual units to deviate from convention, adapt to local conditions and adopt a more ‘trigger-happy’ disposition that proved to be substantially more effective than the more rigid, rule based approach taken by the Dutch forces operating in the same region [50]. Each layer of management is like an abstraction layer. When control is delegated, feedback loops are shortened because layers of abstraction are removed between the controlling entity and the situation at hand. Greater degrees of adaptation are permitted at lower levels of abstraction. This all lends credence to the idea that delegation of control is an important feature of biological systems which is absent from mainstream contemporary artificial intelligence.
3 What is Adaptation?
We will formally establish that delegation of control makes a system more adaptable, all else being equal. Biological systems delegate adaptation down the stack, while computers typically do not. Hence we will argue biological systems are more intelligent, but first we must describe biological systems and computers in comparable terms. We will begin by discussing what it means to learn and adapt, in terms of cause and effect.
@@ 3 What is Adaptation?
To make accurate inferences a system must be able to tell the difference between what it has caused, and what it has merely observed [10, 9, 51]. It is often difficult to determine cause. ‘Confounding’ occurs when an event appears be the cause of the effect we’re trying to predict, but in truth is merely a correlated effect. This can lead one to draw erroneous conclusions about cause and effect.
not the causeeffectcause
@@
By actively intervening in the environment a system can experiment to establish cause and effect. Then, it can construct a causal graph. There are two ways to do that. One option would be to assume the nodes (variables, actions and other abstract objects and properties), and then learn the causal relations between them. This might work well in theory [52], but it requires foreknowledge of objects and properties between which causal relations can be established. The alternative is to assume the causal relations (the edges between nodes in the graph), and then learn the variables (the abstract concepts). This needs only attraction and repulsion from physical states [16, 53], otherwise known as valence. We take this latter approach.
@@
We can frame causal learning in computational terms. Learning nodes in a causal graph is like figuring out which computer program was used to generate a sequence. Assume we have fed a set of inputs into a computer. For each input the computer responded with a particular output. The computer runs a ‘policy’ program which maps inputs to outputs [^8] We’ll call all such programs policies for the sake of this paper. This should not be confused with the more specific notion of a policy from reinforcement learning [19].. That policy is the cause of the input-output history we’ve observed so far. We want to know the policy that caused historical outputs so that we can use it to predict future outputs. The problem is that more than one policy could have produced the history we’ve observed, and each of those policies produce different outputs when we give the computer inputs it hasn’t encountered before. From a naive Bayesian perspective, all of them are equally likely to have been the true underlying cause. If we wish to know which policy actually caused the outputs, we need way to narrow down the possibilities even further.
@@
There are two ways to compare policies that look the same with respect to historical data, but differ with respect to possible future data [46]. One way is to compare their form, and choose programs which are simpler. However, in an interactive setting simplicity is an unreliable indicator of plausibility [47, 46]. The alternative is to compare policies by their function, in terms of how they behave given inputs for which we do not have historical data. Adaptive systems must satisfy functional constraints [54]. We can compare systems by these constraints. Each policy implies a constraint on possible outputs given a set of possible inputs. The weaker that constraint is, the more variety of future behaviour it permits. Previous research showed that if two policies are identical with respect to historical data, then the policy most likely to have generated historical data is the weaker of the two [16]. The policy most likely to have caused behaviour is the weakest, and in experiments a system which preferred weaker policies outperformed one that preferred the simpler policies by $110-500\%$ [55]. This has been used to establish an upper bound on embodied ‘intelligence’ [12]. For the rest of this paper we will use the term ‘weak policy optimisation’ (WPO) to mean a system optimises for weak policies [^9] For a rigorous definition of learning and efficiency, see definition 5 in the appendix.. Included in the appendix of this paper are additional proofs showing WPO is both necessary and sufficient to maximise the probability of ‘generalising’ from one task to another, and thus adaptability.
@@
ML systems do not typically optimise for either weaker constraints on function, or simplicity of form. This at least partially explains their relative sample inefficiency [56, 55]. However, even when ML systems are modified to optimise for weakness, their capacity to do so limited by the “abstraction layer”. An abstraction layer is like a causal isolator [57] between a software “mind” and its environment. For example, hardware like an Apple M2 system on a chip interprets software to physical effect, based on the ARM instruction set architecture. An abstraction layer is just embodiment by another name. Different abstraction layers permit different upper bounds on policy weakness, and thus different upper bounds on embodied ‘intelligence’ [12]. Intelligence is primarily a matter of hardware, not software. In enactivist terms, an intelligent choice of abstraction layer is one that reduces the “big world” of all information to a “small world” of relevant information [58, 59, 60].
@@
Put another way an embodied policy is a physical structure, like a tool. A tool can serve more than one purpose. For example, a hammer is both a weapon and a paperweight. A more specific tool completes fewer possible tasks, while a tool that implies a weaker constraint completes more possible tasks [21]. By preferring weaker policies, one avoids unnecessary assumptions and so adapts as efficiently as possible, meaning the weakest policies complete a ‘wider variety of tasks’ than all others [^10] As an informal epistemological razor, this means “explanations should be no more specific than necessary”. Intuitively, this is a similar idea to the principle of maximum entropy [61, 62], but it accounts for the subjectivity of complexity and the teleological nature of language [46]. For further intuition, a corollary to the idea that a weaker policy is more likely to be the cause of data is the inventors paradox, or equivalently the contravariance principle proposed by Cao and Yamins [63]. They argue tasks that more tightly constrain the space of correct policies make better training sets. A larger, more varied set of inputs and outputs implies a smaller set of correct policies. That makes it easier to identify the policy that generated the inputs and outputs. This supports the claim that weaker policies are more likely to be the cause of behaviour, because a weaker policy is more likely to be a correct policy for tasks that more tightly constrain the set of correct policies. See proofs 1 and 2 for more details.. To complete a ‘wider variety of tasks’ in the embodied sense, a tool must be both more sample and more energy efficient, and so the weakest policy is the most efficient in both senses. Put another way, the same matter can simultaneously contribute to many different tasks or computations superimposed over each other at different scales. A ‘weaker’ policy serves more masters.